### v1

# Diffusion Cocktail: Fused Generation from Diffusion Models

Abstract
Diffusion models excel at generating high-quality images and are easy to extend, making them extremely popular among active users who have created an extensive collection of diffusion models with various styles by fine-tuning base models such as Stable Diffusion. Recent work has focused on uncovering semantic and visual information encoded in various components of a diffusion model, enabling better generation quality and more fine-grained control. However, those methods target improving a single model and overlook the vastly available collection of fine-tuned diffusion models. In this work, we study the combinations of diffusion models. We propose Diffusion Cocktail (Ditail), a training-free method that can accurately transfer content information between two diffusion models. This allows us to perform diverse generations using a set of diffusion models, resulting in novel images that are unlikely to be obtained by a single model alone. We also explore utilizing Ditail for style transfer, with the target style set by a diffusion model instead of an image. Ditail offers a more detailed manipulation of the diffusion generation, thereby enabling the vast community to integrate various styles and contents seamlessly and generate any content of any style.
Figure 1: Diffusion Cocktail demo use cases. Our method primarily focuses on transferring content information between diffusion models. For the top-left example, our method can manipulate the generation process of multiple diffusion models by injecting source image structure, resulting in novel generated images of various styles. We can extend the idea to the other three use cases, including style transfer on real images, stylized regional content injection, and stylized image editing with prompts.
1 Introduction
Figure 2: Comparisons of various diffusion-based image manipulation settings. Unlike all previous methods, Ditail influences the generation of images by incorporating information encoded in multiple DMs. With Ditail, the user can benefit from the largely available set of DMs of dramatically different styles/contents and accurate transfer information between them to get the desirable generation.
Diffusion models (DMs) have shown great success in generating high-quality images [35, 12, 3], conditioning on user inputs such as text snippets [28, 24, 27, 16, 11], reference images [22] or enriched text [6]. A critical feature of the current DMs is the ease of extending and fine-tuning (e.g., using ControlNet [40] or LoRA [13]). The strong capability in image quality and the easy-to-extend feature of DMs result in a vast community of non-industrial/non-academic users fine-tuning the base DMs to generate images of specific styles or specializing certain characters of interest  [^1] To date there are about 7K user-generated DMs checkpoints and about 68K LoRAs available on Civitai: https://civitai.com/.. This gives us access to an extensive collection of fine-tuned DMs with similar model architectures but diverse contents and styles.
Recent work has focused on studying the conceptual/semantic meaning of various components of a DM to interpret its behaviors for controlling the generation process. Jeong et al. [14] show that the essential structure information of the generated images is encoded in the self-attention component in the DM’s U-net vision transformer. Tumanyan et al. [37] utilize such an observation to generate new images based on the reference image and additional text prompts while keeping the spatial information. Kwon and Ye [18] disentangle the content and style information representations in a DM and apply a series of losses to guide an image translation process to manipulate the generated results. Liew et al. [20] mix input semantic concepts by injecting conditioned prompts into the diffusion process and generate images of a blended concept given the prompts.
Even though much work has focused on improving a single DM, rarely have people studied combinations of diffusion models, ignoring the vast collection of DMs created by everyday users. To the best of our knowledge, we are the first to study image generation using multiple DMs and the transfer of information between diffusion models. We propose Diffusion Cocktail (Ditail) that is readily applicable to existing DMs without further training/fine-tuning, thus allowing us to utilize the abundant and fast-growing DM resources efficiently and effectively.
Ditail enables us to accurately transfer the content/structure information between two DMs during the diffusion process. Given two DMs with different styles (e.g., one fine-tuned with watercolor images and the other fine-tuned with line art), in addition to the two images generated by the two DMs individually, Ditail creates two new images, each of which exhibits the style of one DM while complying with the structure of the other DM. This allows users to first generate images using $m$ DMs independently, identify images with desirable contents and styles, and apply Ditail to mix the content and styles for more satisfying results. Ditail can, in principle, generate $\mathcal{O}(m^{2})$ novel images that are not obtainable using a single model from the existing $m$ DMs, which greatly increases the generation flexibility and diversity. DMs with specialized styles or content could get more applications by incorporating other models using Ditail.
Moreover, Ditail is extensible to the style transfer of an input image that is not necessarily generated by a DM. We propose a novel style transfer task: the target style is set according to a generative model, namely a style model, as opposed to a target style image in the classic style transfer setting. Given a reference image and a target style DM, we first extract the content information of the reference image using a general-purpose base DM model (e.g., Stable Diffusion 1.5) and then apply Ditail to transfer style from the base model to the target style model. If no target style model is available, we can fine-tune a DM using a set of style images, thereby essentially composing a mixed target style based on this set of images. At the cost of collecting more images and fine-tuning the model, we benefit from richer information about the target style that no single image could provide.
Finally, Ditail allows more fine-grained control over the process and uncovers the various DM component properties. Ditail could create a smooth interpolation between the content information of the two DMs. We consider both positive and negative prompts in Ditail and provide parameters to control the content corresponding to each prompt, thus creating more diverse generation results. Ditail could also transfer partial content information by applying sub-masks on the self-attention and feature maps in U-Net, empowering users to mix the spatial structures of reference images.
2 Related Work
(a) Overview of the Ditail framework
(b) U-Net Layer
Figure 3: Overview of Ditail for style transfer. The content image first goes through diffusion inversion, ideally using a DM matching the content image’s style. The resulting latents then get injected into the style DM for generation with optional controls such as masking and guidance strength. Given various-style DMs, we generate images with their styles while preserving the content image’s structures. Note if the content image is DM-generated, we may skip the inversion process as we have access to the latent information (novel generation).
Diffusion-based style transfer. Wang et al. [38] first introduce the use of a diffusion model for content extraction and style removal. The diffusion model is fine-tuned to learn disentangled content and style representations. Zhang et al. [41] propose to use text conditioning for style transfer with a diffusion model, where they learn the style of a target image through textual inversion. Yang et al. [39] utilize a pre-trained diffusion model for the DDIM forward step and use the guidance of gradients computed from designed loss terms that preserve the content structure. Everaert et al. [4] take multiple style images and computes high-level statistics to learn a style-specific noise distribution. Instead, Ditail preserves the content/spatial features in a source DM and performs transfer style using a target style DM.
Image manipulation with diffusion models. Pix2pix-zero [25] propose an image-to-image translation method that preserves the source image content without manual prompting to guide the translation. It automatically discovers editing directions that reflect desired edits in the text embedding space and retains the source image content by applying a cross-attention guidance loss throughout the diffusion process.  Hertz et al. [8] identify the importance of cross-attention in controlling the spatial information and its mapping to the text prompt, thus leading to three types of image editing. A few other works have attempted to manipulate image generation with personalized object images as guidance [30, 31, 5]. Orthogonally, our work aims to enable personalized manipulation of image styles through the choice of target style DM, a realistic setup with many fine-tuned DMs available. Tumanyan et al. [37] propose a plug-and-play (PnP) method that edits a given image using a DM with extra prompts to guide the manipulation. PnP first performs diffusion inversion of the given image and extracts features and self-attention maps to record the structure and semantic information. Inspired by PnP, Ditail generalizes the concept of PnP to transfer information between two DMs and facilitates downstream tasks.
Inversion of diffusion models. The inversion problem of a DM aims to find the noise maps as well as the conditioning vector that could generate a given image during the reverse diffusion process. Null-text inversion [23] enables intuitive text-based editing of real images with Stable Diffusion models by using an initial DDIM inversion as an anchor, which only tunes the null-text embedding used in classifier-free guidance. On the other hand, the inversion of the deterministic DDIM sampling process [3] and the inversion of the CLIP [27] have been studied to derive a latent noise map for generating specific real images. Gal et al. [5] propose a textual inversion technique for discovering a new pseudo-word describing the visual concept of a user-provided concept with optimization on only a few images.
Diffusion content-style disentanglement Wang et al. [38] introduce the use of a diffusion model for content extraction and style removal. The diffusion model is then fine-tuned to learn disentangled content and style representations. Jeong et al. [14] show that the structure information is encoded in the self-attention component in the DM’s U-net vision transformer. Based on disentangled representations of content and style, it injects the content information from the h-space (i.e., U-Net’s smallest bottleneck) into the style image through diffusion. Kwon and Ye [18] focuses on image translation through learning disentangled content and style representations. They propose several training objectives to guide the disentangled representation learning with respect to CLIP and [CLS] representations.
3 Method
Preliminaries. Diffusion models [3, 12, 28, 33] are probabilistic generative models, whose generation consists of the forward and the backward process. The forward process is defined as a Markov chain where each step adds a small amount of Gaussian noise to the latent variables. Formally, for a sequence of time steps $t=1,\dots,T$, the process is:
${\boldsymbol{z}}_{t}=\sqrt{\alpha_{t}}{\boldsymbol{z}}_{t-1}+\sqrt{1-\alpha_{t}}{\boldsymbol{\epsilon}}_{t},$ (1)
where ${\boldsymbol{z}}_{t}$ is the latent variable at time step $t$, $\{\alpha_{t}\}$ denotes the variance schedules, $\epsilon_{t}\sim\mathcal{N}(0,I)$. The backward process aims to reconstruct the latent representation from the noise. It is defined as:
${\boldsymbol{\hat{z}}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}({\boldsymbol{\hat{z}}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\boldsymbol{\hat{\epsilon}}}({\boldsymbol{\hat{z}}}_{t},t;\theta)),$ (2)
where ${\boldsymbol{\hat{z}}}_{t-1}$ is the reconstructed latent variable at time $t-1$, $\bar{\alpha}_{t}$ is the cumulative product of $\alpha_{t}$ up to time $t$, ${\boldsymbol{\hat{\epsilon}}}({\boldsymbol{\hat{z}}}_{t},t;\theta)$ is the predicted noise from a model parameterized by $\theta$.
Stable Diffusion [28], a pre-trained text-conditioned Latent Diffusion Model (LDM), has gained attention for its ability to generate high-quality images from textual descriptions. It adopts a pre-trained image autoencoder [17] to perform the diffusion processes in the latent space with an extra condition on the text embedding ${\boldsymbol{c}}$. It uses a U-Net [29] architecture, where each layer consists of a residual block, a self-attention block, and a cross-attention block (Figure 2(b)). At some time step $t$, we denote the output of the $l$-th U-Net block as ${\boldsymbol{z}}_{t}^{l}$; the residual block output as ${\boldsymbol{f}}_{t}^{l}$; the projected self-attention features as ${\boldsymbol{q}}_{t}^{l},{\boldsymbol{k}}_{t}^{l},{\boldsymbol{v}}_{t}^{l}$.
Figure 4: Comparison of content injection. The middle column refers to content injection by replacing ${\boldsymbol{f}}_{t},{\boldsymbol{q}}_{t}$, and ${\boldsymbol{k}}_{t}$. The right column refers to our proposed injection via decoding ${\boldsymbol{z}}_{t}$.
3.1 Ditail Content Injection
Inspired by previous work studying the guidance and translation of DM [14, 38, 37], we transfer the content information from one DM to another by fusing the target DM U-Net hidden states with the ones from the source DM.
[…]
We have two optional control parameters for Ditail. The classifier-free guidance $\omega$ [11] increases the fidelity of the generated image to the positive prompt. The mask $\mathcal{M}$ selects the subset of the latent representations for transfer so that only a portion of the content gets injected. Unlike previous work that injects the features (${\boldsymbol{f}}_{t}$) and self-attention keys and queries (${\boldsymbol{q}}_{t},{\boldsymbol{k}}_{t}$), we inject the content information via decoding the source latent ${\boldsymbol{z}}_{t}$. Comparing the two, injection of ${\boldsymbol{z}}_{t}$ greatly reduces the space required to store the injection information (i.e., from 5.7Gb to 4.7Mb in the case of 50 editing steps). Empirically, we find no significant difference between the two approaches qualitatively (Fig. 4).
Algorithm 1 $\mbox{Ditail}(\{{\boldsymbol{z}}_{t}^{src}\}_{t=1:T},P_{pos},P_{neg};\theta^{tgt})$
1:Input: DM $\theta^{tgt}$, injection latent ${\{{\boldsymbol{z}}_{t}^{src}\}_{t=1:T}}$
2:Optional Control: Classifier-free guidance factor $\omega$, regional injection mask $\mathcal{M}$
3:Output: Target image $\mathcal{I}^{tgt}$
4:${\boldsymbol{e}}_{pos},{\boldsymbol{e}}_{neg}\leftarrow\text{CLIP-enc}(P_{pos}),\text{CLIP-enc}(P_{neg})$
[…]
6:$\tilde{{\boldsymbol{c}}}\leftarrow\text{concat}({\boldsymbol{n}},{\boldsymbol{e}}_{pos},{\boldsymbol{e}}_{neg})$
7:${\boldsymbol{z}}^{tgt}_{T}\leftarrow{\boldsymbol{z}}^{src}_{T}$
8:for $t\leftarrow T\ldots 1$ do
9:    ${\boldsymbol{z}}_{t}\leftarrow\text{concat}({\boldsymbol{z}}^{src}_{t},{\boldsymbol{z}}^{tgt}_{t},{\boldsymbol{z}}^{tgt}_{t})$
10:    ${\boldsymbol{\hat{\epsilon}}}_{t-1}^{pos},{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg}\leftarrow\text{noise-pred-with-inj}\!\left({\boldsymbol{z}}_{t},\tilde{{\boldsymbol{c}}},\mathcal{M},t;\theta^{tgt}\right)$
11:    ${\boldsymbol{\hat{\epsilon}}}_{t-1}\leftarrow{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg}+\omega({\boldsymbol{\hat{\epsilon}}}_{t-1}^{pos}-{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg})$
12:    ${\boldsymbol{z}}^{tgt}_{t-1}\leftarrow\text{DDIM-samp}({\boldsymbol{z}}^{tgt}_{t},{\boldsymbol{\hat{\epsilon}}}_{t-1})$
13:end for
[…]
3.2 Ditail for Novel Generation
The Ditail content injection allows us to generate new images by mixing the representations of two DMs. As described in Alg. 2, given a pair of DMs, we input the positive and negative prompts to generate two images from each DM individually. During the generation process, we can also store the latent representation $\{{\boldsymbol{z}}_{t}^{(1)}\},\{{\boldsymbol{z}}_{t}^{(2)}\}$, and we inject the latent of the first DM to the generation of the second DM and vice versa. By doing so, we generate two more images, each consisting of the structure information of one DM and the style information of the other DM.
Given a set of $m$ DMs, we can use Ditail to generate $m^{2}$ images as opposed to the $m$ images by running every DM independently. This dramatically expands the generation capacity, producing images that are not attainable through a single DM. It also allows the user to have better control over the generation process as well as the integration of more domain-specific DMs. The user may first use some DMs to generate specific content structure and then transfer the spatial information to other DMs with desirable styles. Fig. 5 shows the cross-transfer among three DMs.
3.3 Ditail for Model Style Transfer
As Ditail excels at injecting content information into the diffusion process while preserving the individual DM’s style, we can also treat Ditail as a method for style transfer between models. Essentially, we copy the spatial features and self-attention maps from the content DM and inject those into the style DM. A more precise description for this procedure could be “content transfer” instead of style transfer, as the style DM, which received content information from the other DM, is responsible for generating the target image.
The typical style transfer task takes two images as inputs, and the target is to transfer the style from a style image to the content image while preserving the spatial structure of the content image. Suppose the content image is generated using a DM of the same architecture as the style DM, and we have access to the injection features/self-attention maps; then, we can directly apply style transfer using Ditail as illustrated in the diverse generation. On the contrary, given an arbitrary content image (real or generated), we can first extract the structure information using diffusion inversion and then inject such information using Ditail (Alg. 3).
Algorithm 2 Ditail for Novel Generation
1:Input: Prompts $(P_{pos},P_{neg})$, two DMs $\theta^{(1)},\theta^{(2)}$
2:Output: Two generated images and two novel images
3:$\mathcal{I}^{(1)},\{{\boldsymbol{z}}_{t}^{(1)}\}_{t=1:T}\leftarrow\mbox{DM}(P_{pos},P_{neg};\theta^{(1)})$
4:$\mathcal{I}^{(2)},\{{\boldsymbol{z}}_{t}^{(2)}\}_{t=1:T}\leftarrow\mbox{DM}(P_{pos},P_{neg};\theta^{(2)})$
5:$\mathcal{I}^{(1)\rightarrow(2)}\leftarrow\mbox{Ditail}(\{{\boldsymbol{z}}_{t}^{(1)}\}_{t=1:T},P_{pos},P_{neg};\theta^{(2)})$
6:$\mathcal{I}^{(2)\rightarrow(1)}\leftarrow\mbox{Ditail}(\{{\boldsymbol{z}}_{t}^{(2)}\}_{t=1:T},P_{pos},P_{neg};\theta^{(1)})$
7:return $\mathcal{I}^{(1)},\mathcal{I}^{(2)},\mathcal{I}^{(1)\rightarrow(2)},\mathcal{I}^{(2)\rightarrow(1)}$
Figure 5: Demonstration of style transfer across different DMs. The diagonal images are the non-modified images generated using each diffusion model with the same textual prompt. Every column corresponds to the images generated by one diffusion model, using contents from other models.
Algorithm 3 Ditail for Style Transfer
1:Input: Source content image $\mathcal{I}^{src}$, target style DM $\theta^{tgt}$, inversion DM $\theta^{src}$
2:Optional Control: Condition scaling factors $\alpha,\beta$
3:Output: Target style image $\mathcal{I}^{tgt}$
4:Optionally set the prompts $P_{pos},P_{neg}$ otherwise empty.
5:${\boldsymbol{e}}_{pos},{\boldsymbol{e}}_{neg}\leftarrow\text{CLIP-enc}(P_{pos}),\text{CLIP-enc}(P_{neg})$
6:${\boldsymbol{c}}\leftarrow\alpha\,{\boldsymbol{e}}_{pos}-\beta\,{\boldsymbol{e}}_{neg}$
[…]
8:return $\mbox{Ditail}(\{{\boldsymbol{z}}_{t}^{src}\}_{t=1:T},P_{pos},P_{neg};\theta^{tgt})$
In Line 2-3 of Alg. 3, we first generate a guidance condition vector ${\boldsymbol{c}}$, an interpolation between the positive and negative prompt embedding using two hyper-parameters $\alpha,\beta$. $\alpha$ and $\beta$ are extra knobs that allow the user to control the strength of the guidance for content injection. As shown in Fig 6, a larger $\alpha$ value increases the impact of the source image content, mainly generated by interpreting the positive prompt. Conversely, a larger $\beta$ value does not always yield a significant outcome, as it depends on the semantics of the negative prompt. If the negative prompt is relevant to the content/structure, then $\beta$ will change the resulting content guidance accordingly (Sec. 4 Fig 10). We note that the guidance of $\alpha,\beta$ solely could control the content information: i.e., we use the style DM to generate the output using the guided input latent without injection. But only when combined with the Ditail content injection can we get the best results (more details in Sec. 4).
Figure 6: Visualization of noisy latent after DM inversion with different scaling factor $\alpha$. A larger $\alpha$ leads to a stronger structure prior embedded in the noisy latent, resulting in a more structure-preserved starting point in the reverse sampling process.
We have two extra inputs for applying Ditail on the image style transfer task. Firstly, the user may provide positive and negative prompts to describe the image to help the inversion process. We have also tested empty prompts, prompts generated from image captioning, and irrelevant prompts in Sec. 4 (Fig. 9), which verify that both empty and captioning prompts could generate promising results. Secondly, the user needs to specify a DM for diffusion inversion. Ideally, a DM matching the content image style is desirable. Empirically, the dependence on the inversion model is not strict. Results in Fig. 7 show that the generation results using different inversion DMs are similar for real content images. Compared to the novel generation method as described previously, we may want to apply the style transfer algorithm instead for the following two reasons: 1) the generating DM or the latent representations are not available, and 2) we want to adjust the guidance parameters $\alpha,\beta$ to strengthen or weaken the content transfer.
We note this is a novel style transfer setting as we use DMs to set the styles of interest. This allows us to benefit from the extensive collection of fine-tuned DMs with drastically different styles. If a certain style not included in the available DMs is desirable, we can fine-tune a DM model on images of that style. Compared to the ordinary style transfer, this is equivalent to defining the target style using a set of images instead of only one image. On the one hand, this gives us more flexibility to define the desirable styles, and we may even mix images of different styles into a new style. On the other hand, the set of images tends to provide more stable and comprehensive style information compared to a single image. However, such advantages come at the cost of collecting similar-style images and fine-tuning the DM. But once the target style DM gets trained, we can apply it to transfer styles on numerous content images with Ditail, so the cost goes to zero asymptotically.
Figure 7: Comparison of different inversion models.
4 Experiments
We apply Ditail to multiple tasks and conduct qualitative and quantitative evaluations. Sec. 4.1 performs the proposed novel generation process on generated images and compares the style transfer results with various state-of-the-art methods. Sec. 4.2 ablates the effect of the proposed scaling factors and content injection in terms of fidelity, structure, and style. Sec. 4.3 dives into real image editing empowered by the Ditail framework. More qualitative and quantitative results can be found in the appendix.
Scaling Injection CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
$\alpha=1,\beta=0$ 0.1602 0.0849 173.98
$\alpha=2,\beta=0$ 0.1589 0.0760 173.80
$\alpha=4,\beta=0$ 0.1576 0.0508 188.61
$\alpha=8,\beta=0$ 0.1596 0.0352 214.32
$\alpha=1,\beta=1$ 0.1596 0.0571 186.46
$\alpha=2,\beta=1$ 0.1590 0.0647 179.71
$\alpha=4,\beta=1$ 0.1576 0.0562 182.86
$\alpha=8,\beta=1$ 0.1592 0.0367 212.20
$\alpha=1,\beta=0$checkmark 0.1646 0.0425 196.41
$\alpha=2,\beta=0$checkmark 0.1597 0.0354 194.02
$\alpha=4,\beta=0$checkmark 0.1563 0.0267 209.23
$\alpha=8,\beta=0$checkmark 0.1581 0.0245 228.05
$\alpha=1,\beta=1$checkmark 0.1580 0.0290 204.56
$\alpha=2,\beta=1$checkmark 0.1578 0.0320 197.71
$\alpha=4,\beta=1$checkmark 0.1563 0.0279 203.23
$\alpha=8,\beta=1$checkmark 0.1574 0.0249 226.94
Table 1: Quantitative results for model style transfer on generated images from the GEMRec-18K [7] dataset. We manually pick 5 models with distinctive styles and conduct style transfer on 90 textual prompts. The metrics are averaged over all 20 possible style transfer combinations. The mean CLIP cosine similarity  [9, 26] on source images is 0.1589, and the mean FID [10] distance between source and target images is 227.27. The best-performed metric for each group are shown in bold.
4.1 Results on Generated Images
Dataset and Model Source. Utilizing Ditial for style transfer requires DM checkpoints with various style specializations. The GEMRec-18K [7] dataset contains 18 thousand images generated by 200 text-to-image DMs fine-tuned on Stable Diffusion  [28], where the DM checkpoints were collected from Civitai. Each DM generates 90 images, and the prompts are divided into 12 categories.
Figure 8: Qualitative comparison of style transfer results on generated images drawn from GEMRec-18K [7]. The ControlNet [40] results are produced by its soft edge pre-processing and then fed into a generic img-to-img translation framework.
We manually pick five DM checkpoints with distinctive style specialization and perform any-to-any style transfer using the Ditail framework. We use the images $\{\mathcal{I}^{(1)}_{i}\}_{i=1:90}$ generated by one source DM $\theta^{(1)}$ as the content image and stylize it with another target DM $\theta^{(2)}$, resulting in $\{\mathcal{I}^{(1)\rightarrow(2)}_{i}\}_{i=1:90}$. We compare the qualitative results with other style transfer baselines [21, 40, 15, 41] in Fig. 8. The content and style images are generated by $\theta^{(1)}$ and $\theta^{(2)}$ on the same prompt, respectively. Notably, performing style transfer with Ditail does not require a style image, whereas it is a necessity for conventional style transfer approaches. One key observation here is that the task setting for fine-tuning with one or a few style examples is limited, as it may not secure a stable style learning. Failing to capture certain styles will generally cause a near-unmodified output image, such as the car example for InST [41] and the boy example for the others. It is undeniable that the term “style” has been loosely defined, and most style transfer works tend to specialize in the styles elicited from artistic images. However, we want to expand our scope to a more generalizable setting, which involves practical transfer settings such as “realistic $\rightarrow$ anime” and “watercolor $\rightarrow$ illustration”, or even the nuance style that a single style image or plain text guidance can hardly capture. In such cases, numerous DM checkpoints with stable styles can serve as a plug-and-play choice and exhibit reasonable performance.
4.2 Ablation Study
Quantitative Results. To quantify the change in image fidelity, structure, and style, we use three metrics, respectively. We use the cosine similarity in CLIP-Score [9, 26] to measure the compliance between prompt and image; we use the DINO self-similarity loss [1, 36] to measure the preservation of structure, which essentially compares the self-attention maps between source and target images; we use Fréchet Inception Distance (FID) [10] to the images generated by the target model to measure the closeness of style. We observe the following: First, all CLIP scores do not deviate too much from the source image baseline, and the transferred images from all groups move closer to the distribution of target images. Second, enlarging the scaling factor $\alpha$ tends to degrade the fidelity of generated images, preserve more structure from the source images, and push the transferred images further away from the target distribution. Third, introducing $\beta$ when $\alpha$ is relatively small helps preserve the structure. Forth, the best stylized results are obtained with $\alpha=2$. Lastly, similar to the conclusions from qualitative results, applying content injection alone already provides a strong baseline for structure preservation.
Figure 9: Comparisons of stylized real image editing with various prompts. We set $\omega=7.5$ (left) and $15$ (right) for each image pair.
Figure 10: Ablation on content injection and scaling $(\alpha,\beta)$.
Qualitative Results. As demonstrated in Fig. 10, we can see that solely using scaling factors or content injection can preserve the content structure to some extent. By taking a closer look, we draw the following conclusions: First, scaling $\alpha$ leads to better structure preservation, which complies with the intuition provided by Fig. 6. This plug-and-play technique does not introduce extra computational costs and might be applicable to other tasks that needs structure preservation. Second, introducing $\beta$ is beneficial in polishing the subtle details of the image (e.g., blurriness, lighting and shadow, coloring, etc.) and thereby enhances the overall quality. However, the effect of $\beta$ is quite case-sensitive, as it may not exhibit a quality boost and may potentially cause unnatural coloring of the images, such as the corgi example with $(\alpha,\beta)=(2,2)$. Lastly, applying content injection gives a remarkable structure-preserving performance. Applying both tricks together enables us to keep the content unchanged while shifting to the target style seamlessly.
4.3 Results on Real Images
Scaling Injection CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
$\alpha=1$ 0.0701 0.0920 109.33
$\alpha=2$ 0.0694 0.0852 109.53
$\alpha=4$ 0.0792 0.0609 118.37
$\alpha=8$ 0.0870 0.0478 135.16
$\alpha=1$checkmark 0.0761 0.0517 129.84
$\alpha=2$checkmark 0.0772 0.0449 125.32
$\alpha=4$checkmark 0.0882 0.0322 132.39
$\alpha=8$checkmark 0.0950 0.0297 147.24
Table 2: Quantitative results for model style transfer on 300 real images and captions sampled from the COCO Captions [2] dataset and transfer to 5 stylized DMs. We also use the stylized models to generate pseudo-target images with captions. The mean CLIP cosine similarity [9, 26] on source images is 0.0907, and the mean FID [10] distance between source and target images is 155.43. The best-performed metrics are shown in bold.
Data Source of Real Images. We randomly sample 300 images (with one caption per image) from the COCO Captions [2] dataset. Some demonstration images are drawn from the LAION-5B [32] dataset with textual prompts generated by the captioner model from BLIP [19].
Style Transfer. We perform style transfer on all five stylized models and compare the quantitative results using the aforementioned metrics. The results are presented in Tab. 2. One minor difference here is that we do not have negative prompts from the collected data, plus the benefit of introducing $\beta$ is not generalizable enough according to Tab. 1, so we skip the comparison on $\beta$ for this section. Overall, the quantitative patterns remain largely the same compared to the results on generated images, except that the best-performed CLIP cosine similarity is dominated by the groups with the largest $\alpha$. We hypothesize that this is because the caption of real images does not necessarily capture all the components of the images, whereas the generated images are produced from the given prompts. With a larger $\alpha$, the CLIP scores may still fluctuate for generated images but increase in the case of real images.
Stylized Image Editing. The Ditail framework enables stylized real-world image editing as well. Some qualitative results over various prompt types are shown in Fig. 9. The patterns we observed are as follows: First, null prompts hardly exhibit editing power and generally result in slightly stylized output. Second, a larger guidance scale $\omega$ [11] allows the target models to be more creative. Third, given that we enforce a strong structure emphasis through condition scaling and content injection, the editing prompts that change the number of objects usually give unsatisfactory results, whereas the semantic class of the objects might be modified effectively, even for completely unrelated prompts (e.g., car $\rightarrow$ boar, man $\rightarrow$ chicken, etc.). Lastly, the editing prompts that modify the colors (e.g., blue car to white car) can be quite effective in practice.
5 Conclusion
We propose Ditail, a training-free method to transfer content and style information between DMs. Ditail allows for novel generation by incorporating various DMs, giving users more control and flexibility and benefiting from the extensive collection of DMs. Ditail applies to style transfer tasks with the target style defined by a DM instead of images and has the promise of fine-grained manipulations such as prompt guidance strength and partial transfer for controllable image generations. Ditail is also generalizable to stylized image editing on real images. Our future work involves integrating Ditail to other emerging methods and architectures for better efficiency and controllability.
Acknowledgments
This work is supported in part by the Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning at NYU Shanghai, NYU Shanghai Boost Fund, and NYU High Performance Computing resources and services.
Supplementary Material
6 Model Checkpoints Information
We manually select 5 DM checkpoints with distinctive style specialization from civitai  [^2] https://civitai.com. Fig. 11 lists the important information about these models, including the style keyword we assign and the tags provided by model uploaders. We also select some LoRA checkpoints from civitai and liblib  [^3] https://www.liblib.ai, whose information could be found in Fig. 12.
Figure 11: List of selected DM checkpoints from Civitai. Style keyword for each checkpoint are assigned by us. Example images are drawn from GEMRec-18K [7].
Figure 12: Partial list of selected LoRA checkpoints from civitai and liblib. Style keyword for each checkpoint are assigned by us. Example images are generated by LoRA uploaders.
7 Implementation Details
We implement the Ditail framework based on PyTorch and Diffusers  [^4] https://huggingface.co/docs/diffusers/index. Following prior works [37], we use Denoising Diffusion Implicit Models (DDIM) scheduler [34] with 1000 inversion steps and 50 sampling steps. Note that the inversion steps can be further reduced without a significant loss of output quality. In the project demo, we use 50 steps for both inversion and sampling and set $(\alpha,\beta)=(2.0,0.5)$ by default. With a Nvidia A10G-small GPU on HuggingFace, processing a single image takes roughly 15 seconds using 16-bit floating point precision.
8 Pseudocode for noise-pred-with-inj($\cdot$)
1# Set injection threshold
2total_steps = 50
3thresh_res = 0.8 * total_steps
4thresh_attn = 0.5 * total_steps
5
6
7# Input:
8#   - m: ResNet block to be forwarded
9#   - mask: Optional mask for local injection
10#   - x: Input hidden states
11# Output:
12#   - Forwarded hidden states
13def inj_forward_res(m, mask, x):
14    h_src, h_pos, h_neg = m.forward(x).chunk(3)
15    h_pos = mask * h_src + (1 - mask) * h_pos
16    h_neg = mask * h_src + (1 - mask) * h_neg
17    return concat(h_src, h_pos, h_neg)
18
19
20# Input:
21#   - m: Self-attention block to be forwarded
22#   - mask: Optional mask for local injection
23#   - x: Input hidden states
24# Output:
25#   - Forwarded hidden states
26def inj_forward_attn(m, mask, x):
27    q, k, v = m.to_q(x), m.to_k(x), m.to_v(x)
28    q_src, q_pos, q_neg = q.chunk(3)
29    k_src, k_pos, k_neg = k.chunk(3)
30    q_pos = mask * q_src + (1 - mask) * q_pos
31    q_neg = mask * q_src + (1 - mask) * q_neg
32    k_pos = mask * k_src + (1 - mask) * k_pos
33    k_neg = mask * k_src + (1 - mask) * k_neg
34    q = concat(q_src, q_pos, q_neg)
35    k = concat(k_src, k_pos, k_neg)
36    return m.attn_forward(q, k, v)
37
38
39# Input:
40#   - unet: U-Net model
41#   - x: Input noisy latents
42#   - t: Current time step
43#   - inj_ids: List of block IDs with injection
44# Output:
45#   - Denoised latent
46def noise_pred_with_inj(unet, x, t, inj_ids):
47    for i, m in enumerate(unet.blocks()):
48        if i in inj_ids:
49            if is_res(m) and t > thresh_res:
50                m.forward = inj_forward_res
51            if is_attn(m) and t > thresh_attn:
52                m.forward = inj_forward_attn
53    return unet.forward(latents)
Listing 1: Pseudocode for noise-pred-with-inj
9 Inversion Model Comparison
One natural variant of Ditail style transfer is to invert a source image with non-original DMs. In this section, we perform both qualitative and quantitative investigations on the pros and cons of such a variant. As shown in Tab. 3, performing inversion with the target DM enhances structure preservation but degrades image fidelity and style. Some qualitative results are presented in Fig. 13. The first row demonstrates a failure case of inverting with source DM. This happens rarely when the source DM possesses a strong style, yet the target DM fails to denoise the strongly stylized latents properly and causes a collapse of content and structure. In such cases, inverting with target DMs can give better results. However, this also inevitably degrades the stylizing capability. As shown in the second and third rows, though inverting with the target DM ensures better structure preservation, this also leads to less stylized output. For real images, the quantitative results in Tab. 4 show similar patterns. Note that there is no source DM for real-world images, but its distribution can be best simulated by photo-realistic DMs, allowing more flexibility in the structure and thus resulting in more stylized output.
As an additional reference, we also attach the noisy latents inverted with different DMs in Fig. 14. We can observe that the noisy latents do exhibit some style-specific patterns. For example, the latents inverted with the Abstract DM are less colorful than the others. This is actually opposed to the patterns we observed from the output images, as inverting with the target DM does make the latents more “stylized”. Nevertheless, we hypothesize that target DMs tend to exhibit less capability of modifying the structure produced by other DMs, which may potentially lead to a better-preserved structure and a weakened style.
Inversion Model Scaling CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
Source $\alpha=1,\beta=0$ 0.1646 0.0425 196.41
$\alpha=2,\beta=0$ 0.1597 0.0354 194.02
$\alpha=4,\beta=0$ 0.1563 0.0267 209.23
$\alpha=8,\beta=0$ 0.1581 0.0245 228.05
Target $\alpha=1,\beta=0$ 0.1601 0.0288 204.87
$\alpha=2,\beta=0$ 0.1574 0.0262 202.09
$\alpha=4,\beta=0$ 0.1545 0.0236 212.77
$\alpha=8,\beta=0$ 0.1555 0.0232 230.18
Table 3: Quantitative comparison of using different DMs for inversion on generated images. The mean CLIP cosine similarity  [9, 26] on source images is 0.1589, and the mean FID [10] distance between source and target images is 227.27. The best-performed metric for each group is shown in bold.
Inversion Model Scaling CLIP $\uparrow$ DINO $\downarrow$ $\text{FID}_{\text{tgt}}\downarrow$
Realistic $\alpha=1$ 0.0761 0.0517 129.84
$\alpha=2$ 0.0772 0.0449 125.32
$\alpha=4$ 0.0882 0.0322 132.39
$\alpha=8$ 0.0950 0.0297 147.24
Target $\alpha=1$ 0.0750 0.0372 153.06
$\alpha=2$ 0.0781 0.0324 154.73
$\alpha=4$ 0.0872 0.0283 157.61
$\alpha=8$ 0.0952 0.0282 171.84
Table 4: Quantitative comparison of using different DMs for inversion on real images. The mean CLIP cosine similarity  [9, 26] on source images is 0.0907, and the mean FID [10] distance between source and target images is 155.43. The best-performed metrics are shown in bold.
10 More Qualitative Results
We attach more qualitative results produced with the Ditail framework in this section. Fig. 15 presents the matrix view of cross-transfer across DMs of various styles (as in Fig. 5). Fig. 16 and Fig. 17 shows real image style transfer results with different model configurations and scaling factors. Fig. 18 ablates content injection and condition scaling on real images. Fig. 19 investigates the reversibility of Ditail style transfer. Fig. 20 explores the stability of the reconstructed content by performing Ditail multiple times. Please find the takeaways in the captions of each figure.
Figure 13: Comparison of performing inversion with source and target DMs.
Figure 14: Visualizations of the noisy latents inverted by different DMs. We use the same source image as in Fig. 6.
Figure 15: Demonstration of style transfer across various DMs. Similar to Fig. 5, images on the diagonal are those generated by the original DMs, whereas the others are the novel results. Ditail enables us to mix content and style information easily on images of all types.
Figure 16: More style transfer results (DM only) on real images, which demonstrates Ditail’s high generalizability.
Figure 17: More style transfer results (DM+LoRA) on real images with different scaling factors. During generation, trigger words for each LoRA are appended to the prompt, and we set the scale of LoRA weight to be 0.7, as recommended by LoRA uploaders. In most cases, we observe evident transformations in image style and content, while the structure of the source image is largely preserved. We also notice that the effect of the same LoRA on different checkpoints varies, which entails the DM-LoRA compatability problem.
Figure 18: Ablation studies on content injection and scaling factors for real images. We can see that applying either of them helps preserve the structure; otherwise, the output images will deviate from the original source content (i.e., close to vanilla text-to-image generation).
Figure 19: Results of transferring the novel images back to their original style. In general, the content deviation depends on 1) the content complexity and 2) the style gap between the source and target DMs. For example, the duckweed (complex) is subject to change after the two-step transfer; whereas the car (simple) remains largely the same. It is hard to restore the lighting as well (room example).
Figure 20: Results of reconstructing the input images with a single DM multiple times. In general, a larger $\alpha$ helps stabilize the content in the source image. However, it may lead to a weird coloring issue when the magnitude of the condition vector is too large.